ProbabilityMedium
Question
Two events A and B have probabilities 0.25 and 0.50 respectively. The probabilitity that both A and B occur simultaneoushly is 0.14 Then, the probability that neither A nor B occurs is
Options
A.0.39
B.0.25
C.0.11
D.None of these
Solution
Given, P(A) = 0.25, P(B) = 0.50, P(A ∩ B) = 0.14
∴ P(A ∪ B) = P(A) + P(B) - P(A ∩ B)
⇒ P(A ∪ B) = 0.25 + 0.50 - 0.14
P(A ∪ B) = 0.61

∴ P(A ∪ B) = 1 - P(A ∩ B) = 1 - 0.61 = 0.39
∴ P(A ∪ B) = P(A) + P(B) - P(A ∩ B)
⇒ P(A ∪ B) = 0.25 + 0.50 - 0.14
P(A ∪ B) = 0.61

∴ P(A ∪ B) = 1 - P(A ∩ B) = 1 - 0.61 = 0.39
Create a free account to view solution
View Solution FreeMore Probability Questions
A tetrahedral fair dice whose faces are numbered $1,2,3$ and 4 . The die is thrown and the face on which the pyramid lan...The probability of India winning a test match against West Indies is $\frac{1}{2}$. Assuming independence from match to ...The probability that a non leap year will have 52 Fridays is-...If p1 and p2 are the probabilities of two independent events then (1 − p1 − p2 + p1p2) is the probability of...Pal′s gardner is not dependable, the probability that he will forgot to water the rose bush is 2/3. The rose bush ...